# SignBot: Learning Human-to-Humanoid Sign Language Interaction

### v1

@@ 5.3 Naturalness: Sim-to-Real Human-Robot Interaction
Limitation. 1) Limitations in Sign Language Translation. Current state-of-the-art open-source models for sign language translation still exhibit errors, making it difficult to handle complex expressions. Although not the primary focus of this work, more advanced models can be seamlessly integrated into our framework in the future. 2) Latency in Human-Robot Interaction. The robot’s reasoning module currently relies on separate models for dialogue, translation, and sign language generation, resulting in latency. Unifying these components into a vision-language-action model that directly produces robot-adapted sign language actions could greatly improve real-time performance.
6 Conclusion
@@
We introduce SignBot, a human-robot sign language interaction framework for humanoid robots that incorporates an embodied cerebellum + cerebral cooperation mechanism. This framework has been validated across various sign language motions, demonstrating exceptional accuracy, generalization, naturalness, and adaptability across diverse sign language scenarios. In particular, the cerebellum + cerebral cooperation mechanism in SignBot achieves reliable performance in daily communication through the sign language translator, response, and generator. We think SignBot is a foundational solution for sign language applications, such as daily sign language robots serving the DHH community. A promising direction for future research is integrating SignBot with manipulation skills. In this way, we can not only communicate with the DHH community but also assist them in completing tasks.
@@ B.1. Task, Regularization and Penalty Rewards
DoF pos limit $1(\mathbf{q}_{t}\notin[\mathbf{q}_{\min},\mathbf{q}_{\max}])$ $-1e^{-2}$
Alive $1$ $1$
@@
Drag $\sum_{i}||\mathbf{v}_{foot}||*\mathrm{1}^{new\,contact}$0.1
Contact Force $1\{|F^{i}_{t}|\geq F_{th}\}*(|F^{i}_{t}|-F_{th})$3e-3
Stumble $1\{\exists i,|F^{i}_{xy}|>4|F^{i}_{t}|\}$2.0
DoF Acceleration $|\dot{\mathbf{q}}|^{2}$3e-7
Action Rate $|\mathbf{a}_{r}-\mathbf{a}_{r}|$1e-1
Energy $|\dot{\mathbf{q}}|^{2}$1e-3
DoF Limit Violation $1\{\mathbf{q}_{i}>\mathbf{q}_{max}||\mathbf{q}_{i}<\mathbf{q}_{min}\}$10.0
DoF Deviation $||\mathbf{q}_{default}^{low}-\mathbf{q}^{low}||^{2}$1e-1
Vertical Linear Velocity $\mathbf{v}_{2}$1.0
Horizontal Angular Velocity $|\mathbf{\omega}_{xy}|^{2}$0.4
Projected Gravity $\|\mathbf{g}_{xy}\|^{2}$2.0
Table 3: Task, Penalty and regularization reward expressions for different terms.
@@ B.2. PPO Parameters
We apply PPO to train the H1 robot. Table 4 shows the PPO parameters.
Table 4: PPO Hyperparameters
@@
Figure 8: Sign Language Alignment between Human and Legged Robot.
@@ B.3. Domain Randomization
Domain randomization is implemented across all training environments. This technique introduces variability in various environmental parameters, including friction and gravity. Detailed specifications of the domain randomization can be found in Table 5.
@@
Table 5: Domain randomization parameters: we include the friction, gravity, motor strength, base CoM/mass offset of torso link, link mass, PD Gains and an external perturbation.
@@ B.4. Running Setting
To eliminate redundancy caused by the same sign language data being demonstrated by different users, we remove duplicate samples and select the 6576 sign language data for training baselines. We run experiments with three different seeds and present the mean ± std results for each algorithm. To ensure a fair comparison, we maintain the same settings for all comparison baselines. We use 123, 321, and 1 as the experimental seeds, and we present the mean ± standard deviation (std) for each evaluated algorithm. The training process for the motion control policy utilizes 12GB of GPU memory across 1 RTX 4090 and typically runs around 48 hours.
@@ B.5. Torque Setting
The PD gains, characterized by stiffness and damping values, used in the IssacGym simulator are detailed in Table 6.
Table 6: Torque Parameters
@@ C. Case Study
We conduct a case study (see Figure 7), and it can be observed that while whole-body tracking provides relatively stable tracking of the lower body, the upper body exists some errors due to its inability to keep up with the motion frequency. Additionally, if SignBot does not track the posture of the lower body, it becomes very difficult to maintain stability. In addition, because our H1 robot has a pair of larger dexterous hands and a pair of modified wrists, the weight of the upper body has increased. Employing upper body tracking/whole body tracking methods makes it difficult for the robot to maintain its center of gravity while standing still. Therefore, we decouple the policy to allow RL to primarily learn to track the lower body, reducing the difficulty of training.
Figure 8 shows the interaction between human and H1 robot.
D. Sign Language Interaction Experiment Supplement
D.1. Prompt
The sign language responder needs a suitable prompt template, which mainly depends on your scenario. We briefly provide a template for family life, and you can also encourage it to ask you questions or interact. In experiments, it may sometimes produce some strange responses, so it is occasionally necessary to guide the LLM and set word limits.
Please act as a family member in daily life, such as my siblings. The topics can include greetings, family, work, food, travel, etc. Responses should be concise, reasonable, and logical according to everyday life. You can interact appropriately by asking reasonable questions or providing sensible answers based on what I say, and the responses should not be lengthy. Your response should think and speculate about the meaning of what the other person is saying as much like a human as possible, and answer using everyday logic. For example, if someone invites you to run together and you want to join, you need to have some reasonable basis, such as you enjoy running or you believe running is beneficial for your health. If you don’t want to run with them, it might be because you’re not feeling well, or you have other important matters to attend to, such as studying, or due to external reasons like bad weather today.
D.2 More Expressive Interaction Examples
Please check Figure 8. Even without the H1 hardware, we can attempt sign language interaction tasks in simulation.

### v3

@@ V-C Naturalness: Sim-to-Real Human-Robot Interaction
Translation Method BLEU-1 $\uparrow$ BLEU-4 $\uparrow$ ROUGE $\uparrow$ Generation Method DTW-PA-JPE (body) $\downarrow$ DTW-PA-JPE (hand) $\downarrow$
C2RL (CSL) [23] 49.32 21.61 48.21 NSA (ASL) [13] 7.83 7.33
Translator (ASL) 40.20 14.90 36.01 Generator (ASL) 6.82 2.35
Translator (CSL) 55.08 26.36 56.51 Generator (CSL) 6.24 1.71
Translator (CSL, Deployment) 53.20 20.53 54.48 Generator (CSL, Deployment) 7.63 2.19
TABLE III: Evaluation of Sign Language Translation and Generation Models in the simulation and real environment.
We evaluate the performance of the sign language translation and generation modules in interaction. We compare our modules with the current SOTA baselines on the ASL and CSL datasets using the test sets (since the translator is a gloss-free method, we compare it with the gloss-free baseline instead of the gloss-based baseline). Additionally, we test the effectiveness of both modules using real-world sign language videos and dialogue corpus. Since our sign language users are familiar with CSL, the actual performance is only evaluated on CSL. Table III shows that our modules outperform the existing SOTA baseline. "Deployment" refers to the evaluation using our real test data from non-CSL test sets during the actual device deployment phase. Furthermore, since the dialogue content mainly falls within the vocabulary and scope covered by the CSL dataset, the performance degradation is not significant, providing a basic guarantee for sign language interaction.
V-D Interactivity: Human Feedback from the DHH Community
We conduct a user study and invite members of the DHH community to evaluate this work. Due to constraints of the experimental venue and hardware safety concerns, we do not involve DHH individuals in actual interactions. As an alternative, we provide comprehensive documentation and video recordings of the entire sign language interaction process to facilitate their assessment.
Evaluator Satisfaction Enjoyment Naturalness Accuracy Understanding Convenience
Evaluator 1 6 5 6 5 6 8
Evaluator 2 7 8 7 6 5 5
Evaluator 3 8 7 9 7 7 8
Evaluator 4 7 6 7 7 7 8
Evaluator 5 6 6 6 5 6 8
Evaluator 6 6 9 5 7 6 9
Evaluator 7 9 9 7 8 6 5
Evaluator 8 3 3 3 3 3 3
Evaluator 9 1 1 1 1 5 6
Evaluator 10 8 6 7 7 7 6
TABLE IV: Evaluation Results from the DHH Community
Experiment Setup: Among these people, there are young and elderly individuals, all coming from different places of birth. To ensure fairness, we do not collect any personal or private information from the experiment participants. The assessment covers the following aspects: 1) Overall satisfaction with the system, 2) enjoyment of the interactive experience, 3) naturalness and fluency of responses, 4) accuracy of motion execution, 5) ability to understand human-expressed intentions, and 6) whether the system could bring convenience to their daily lives. Ratings are given on a scale of 1–10, where 10 represents the highest score. A score of 6 represents a baseline level of satisfaction.
Table IV presents the individual ratings from all 10 evaluators. Most evaluators give scores above 6 in the majority of categories, indicating general satisfaction with the system’s performance. The convenience criterion received the highest average score (mean: 6.60), suggesting that evaluators think this research is meaningful and has the potential to positively impact their daily lives. If we exclude Evaluators 8 and 9 (likely outliers due to unfamiliar regional signs), the average score across all criteria increases significantly (e.g., Satisfaction average rises from 6.10 to 7.13). Naturalness (from 5.80 to 6.75) and Accuracy (from 5.60 to 6.50) receive moderately high scores, despite the absence of facial expressions or lip movements.
Limitation. Based on feedback from members of the DHH community, we have summarized several current limitations of our framework: 1) Limitations in Sign Language Translation and Generation. Current state-of-the-art open-source models for sign language translation and generation still exhibit errors. Although we currently use public sign language datasets, our research has found that sign language can vary across different regions within the same country. 2) Limitations of hardware support. Our robot cannot support facial expressions or lip movements, which may affect the understanding of certain sign language content.
Despite these limitations, a significant portion of the DHH community strongly affirms both the technical quality of our implementation and the social value of our research.
VI Conclusion
@@
We introduce SignBot, a human-robot sign language interaction framework that incorporates an embodied cerebellum + cerebral cooperation mechanism. This framework has been validated across various sign language motions, demonstrating exceptional accuracy, generalization, naturalness, and adaptability across diverse sign language scenarios. In particular, the cerebellum + cerebral cooperation mechanism in SignBot achieves reliable performance in daily communication through the sign language translator, response, and generator. We think SignBot is a foundational solution for sign language applications, such as daily sign language robots serving the DHH community. In the future, we will focus on enabling the robot to exhibit facial expressions, enhancing the interaction experience, and collecting field data in different provinces to accommodate regional variations in sign language.
APPENDIX
@@ B.1. Task, Regularization and Penalty Rewards
Table V showcases the Task, Regularization, and Penalty Rewards in RL training.
Task rewards $r_{t}$ assess how effectively the agents achieve the goals of the current task, such as approaching the target joint Dof pos. These rewards are closely correlated with the optimality of the current control policy, making them an appropriate objective to maximize in this context. Penalty rewards $r_{p}$ aim to prevent undesirable events. For example, these penalties will take effect when the humanoid falls or the robot violates dynamic constraints, including torque limits, joint position limits, and so on. Regularization rewards $r_{r}$ are used to align the behavior of the humanoid with human preferences regarding motion styles, safe deployment, and other kinematic or dynamic concerns.
@@
DoF pos limit $1(\mathbf{q}_{t}\notin[\mathbf{q}_{\min},\mathbf{q}_{\max}])$ $1e^{-2}$
Alive 11 11
@@
[…]
Contact Force $1\{|F^{i}_{t}|\geq F_{th}\}*(|F^{i}_{t}|-F_{th})$ -3e-3
Stumble $1\{\exists i,|F^{i}_{xy}|>4|F^{i}_{t}|\}$ -2.0
DoF Acceleration $|\dot{\mathbf{q}}|^{2}$ -3e-7
Action Rate $|\mathbf{a}_{r}-\mathbf{a}_{r}|$ -1e-1
Energy $|\dot{\mathbf{q}}|^{2}$ -1e-3
DoF Limit Violation $1\{\mathbf{q}_{i}>\mathbf{q}_{max}||\mathbf{q}_{i}<\mathbf{q}_{min}\}$ -10.0
DoF Deviation $\mathbf{q}_{default}^{low}-\mathbf{q}^{low}||^{2} -$1e-1
Vertical Linear Velocity $\mathbf{v}_{2}$ -1.0
Horizontal Angular Velocity $|\mathbf{\omega}_{xy}|^{2}$ -0.4
Projected Gravity $\|\mathbf{g}_{xy}\|^{2}$ -2.0
TABLE V: Task, Penalty and regularization reward terms.
@@ B.2. PPO Parameters
Table VI shows the PPO parameters.
In our PPO implementation, the actor network is composed of a convolutional neural network (CNN) followed by multilayer perceptrons (MLPs), where the CNN is primarily responsible for extracting features from the historical observation inputs. The critic network, in contrast, is constructed purely from MLPs, leveraging the extracted state features for value estimation.
TABLE VI: PPO Hyperparameters
@@ B.3. Domain Randomization
Domain randomization is implemented across all training environments. This technique introduces variability in various environmental parameters, including friction and gravity. Detailed specifications of the domain randomization can be found in Table VII.
@@
TABLE VII: Domain randomization parameters: we include the friction, gravity, motor strength, base CoM/mass offset of torso link, link mass, PD Gains, and an external perturbation.
@@ B.4. Running Setting
We run experiments with three different seeds (123, 321, 1) and present the mean ± std results for each algorithm. The training process for the motion control policy utilizes 12GB of GPU memory across 1 RTX 4090 and typically runs around 48 hours.
To ensure a fair comparison, we maintain the same settings for all comparison baselines.
@@ B.5. Torque Setting
The PD gains, characterized by stiffness and damping values, used in the IssacGym simulator are detailed in Table VIII.
TABLE VIII: Torque Parameters
@@ C. Case Study
We conduct a case study (see Figure 7), and it can be observed that while whole-body tracking provides relatively stable tracking of the lower body, the upper body exists some errors due to its inability to keep up with the motion frequency.
Additionally, if SignBot does not track the posture of the lower body, it becomes very difficult to maintain stability. In addition, because our H1 robot has a pair of larger dexterous hands and a pair of modified wrists, the weight of the upper body has increased. Employing upper body tracking/whole body tracking methods makes it difficult for the robot to maintain its center of gravity while standing still. Therefore, we decouple the policy to allow RL to primarily learn to track the lower body, reducing the difficulty of training.